# GPT4RoI: Instruction Tuning Large Language Model on Region-of-Interest

### v4

@@ 5.2 Region Recognization
Open-Vocabulary Recognition The results are obtained by calculating the semantic similarity between the generated region caption of GPT4RoI and the vocabulary lists of each dataset (as the study Yuan et al. (2023)) and then selecting the category with the highest similarity as the final result. The input to the CLIP-Surgery-ViT-L is the cropped region with a size of 512$\times$512. As shown in Table 4, constrained by the resolution of 224$\times$224, GPT4RoI exhibits slightly lower performance than CLIP-Surgery-ViT-L only at the instance level. Considering all metrics score on two datasets, GPT4RoI demonstrates robust and comprehensive region recognition capabilities compared to our concurrent works that employ alternative reference formats.
Method Cityscapes ADE20K-150
PQ AP mIoU PQ AP mIoU
CLIP-Surgery-ViT-L 27.24 28.35 21.92 26.55 29.70 21.42
Kosmos-2 12.09 9.81 13.71 6.53 4.33 5.40
Shikra-7B 17.80 11.53 17.77 27.52 20.35 18.24
GPT4RoI-7B 34.70 21.93 36.73 36.32 26.08 25.82
Table 4: Recognition performance on panoptic segmentation (PQ), instance segmentation (AP) and semantic segmentation (mIoU) upon the validation sets of Cityscapes Cordts et al. (2016) and ADE20K Zhou et al. (2017). means concurrent work.
Model BLEU@4 METEOR ROUGE CIDEr
GRiT - 17.1 - 142.0
Shikra-7B 8.9 15.2 30.3 115.8
GPT4RoI-7B 10.5 16.5 33.4 134.5
GPT4RoI-7B 11.5 17.4 35.0 145.2
GPT4RoI-13B 11.7 17.6 35.2 146.8
Model Code #Params Val Acc.(%) Test Acc.(%)
ViLBERT Lu et al. (2019) Y 221M 72.4 74.5 54.0 73.3 74.6 54.8
Unicoder-VL Li et al. (2019a) Y - 72.6 74.5 54.5 73.4 74.4 54.9
VLBERT-L Su et al. (2019) Y 383M 75.5 77.9 58.9 75.8 78.4 59.7
UNITER-L Chen et al. (2020) Y 303M - - - 77.3 80.8 62.8
ERNIE-ViL-L Yu et al. (2021) Y - 78.52 83.37 65.81 79.2 83.5 66.3
MERLOT Zellers et al. (2021) Y 223M - - - 80.6 80.4 65.1
VILLA-L Gan et al. (2020) Y - 78.45 82.57 65.18 78.9 82.8 65.7
RESERVE-L Zellers et al. (2022) Y 644M - - - 84.0 84.9 72.0
VQA-GNN-L Wang et al. (2022) Y 1B+ - - - 85.2 86.6 74.0
GPT4RoI-7B (ours) Y 7B+ 87.4 89.6 78.6 - - -
VLUA+@Kuaishou N - - - - 84.8 87.0 74.0
KS-MGSR@KDDI Research and SNAP N - - - - 85.3 86.9 74.3
SP-VCR@Shopee N - - - - 83.6 88.6 74.4
HunYuan-VCR@Tencent N - - - - 85.8 88.0 75.6
Human Performance Zellers et al. (2019a) - - - - - 91.0 93.0 85.0
GPT4RoI-13B (ours) Y 13B+ - - - 89.4 91.0 81.6
Region Caption We report the scores of BLEU, METEOR, ROUGE, and CIDEr for different methods on the validation set of Visual Genome Krishna et al. (2017). As shown in Table 5, compared to the concurrent work Shikra Chen et al. (2023b) with textual coordinates as region reference, GPT4RoI shows clearly better region caption ability. After fine-tuning, GPT4RoI outperforms the previous state-of-the-art specialist model GRiT Wu et al. (2022) by a significant margin, without any additional techniques or tricks. Additionally, we observe that the performance of GPT4RoI-7B and GPT4RoI-13B is comparable, suggesting that the bottleneck in performance lies in the design of the visual module and the availability of region-text pair data. These areas can be explored further in future work.
5.3 Region Reasoning
Visual-7W Zhu et al. (2016) and Visual Commonsense Reasoning Zellers et al. (2019a) employ multiple-choice settings that require both recognition ability and commonsense reasoning to select the correct choice. Calculating the similarity of LLM’s output with the correct choice is quite a challenge. This is because it is hard to find a model that can capture the reasoning similarity. Therefore, we finetune GPT4RoI to these two datasets to align with the answer format, following conventional methods.
Visual-7W Visual-7W Zhu et al. (2016) is a PointQA dataset that contains a which box setting. Here, the model is required to choose the appropriate box among four options, based on a given description. For example, a question might ask, “Which is the black machine under the sign?”. This type of question not only tests the model’s object recognition but also its ability to determine the relationship between objects.
To prevent information leakage, we remove overlapping images with the test set from Visual Genome Krishna et al. (2017). As shown in Table 7, the results demonstrate that the 13B model outperforms the 7B model by a significant margin. This finding suggests that the reasoning ability heavily relies on the LLM.
Model LSTM-Att CMNs 12in1 GPT4RoI-7B GPT4RoI-13B
Acc(%) 56.10 72.53 83.35 81.83 84.82
Table 7: Accuracy on Visual-7W. GPT4RoI-13B achieves SoTA performance.
Visual Commonsense Reasoning Visual Commonsense Reasoning (VCR) offers a highly demanding scenario that necessitates advanced reasoning abilities, heavily relying on common sense. Given the question(Q), the model’s task is not only to select the correct answer(A) but also to select a rationale(R) that explains why the chosen answer is true. We give a more detailed explanation of each metric in our Appendix A.3.
6 Conclusions
7 Limitations & Social Impact
7.1 Limitations
In our exploration, we find GPT4RoI produces failure cases as shown in Section. A.5. The temporary limitation of GPT4RoI falls in instruction obfuscation and misidentification of fine-grained information within the region. We summarise them into 3 items to point out the limitations and potential solutions
  • Model architecture. We find that 224 $\times$224 input image resolution struggles with understanding smaller regions. However, if we switch to a larger resolution, we must consider the potential burden on inference speed from global attention ViT architecture, while the more efficient CNN architecture or sliding window attention has no available pre-trained large-scale vision encoder like CLIP ViT-L/14.
  • Region-level instructions. Although we have generated instructions for each task from existing open-source datasets, users in practical applications may ask various questions about an arbitrary number of regions, and the existing data may not contain satisfactory answers.
  • Interaction mode. Currently, GPT4RoI only supports natural language and bounding box interaction. Incorporating more open-ended interaction modes such as point, scribble, or image-based search could further improve the user interaction experience.
7.2 Social Impact
The ability of GPT4RoI to analyze and interpret visual content raises privacy concerns. If misused or unregulated, it could compromise individuals’ privacy by extracting sensitive information from images without their consent. The widespread implementation of such strong image-understanding algorithms may automate jobs that traditionally rely on human interpretation of visual data. This could lead to unemployment or job displacement for individuals working in fields such as image recognition, analysis, or interpretation. It is crucial to emphasize that the social impact of any technology depends on its development, deployment, and regulation. Ethical considerations, responsible use, and transparency in algorithmic decision-making can help mitigate the potential negative effects and maximize the positive contributions of GPT4RoI in society.
Appendix A Appendix
Figure 4: A more detailed framework of GPT4RoI.
In this appendix, we provide a detailed method architecture figure. We then discuss training-related details, including hyperparameters and instruction templates used in each stage and task. Specifically, we give an introduction to VCR dataset and describe how we utilize the VCR dataset. We also compare the design philosophy with methods using textual coordinates in LLM and provide a brief overview of concurrent works. Finally, we analyze some error cases and propose potential improvements for future exploration.
A.1 Detailed Architecture
Details of our region feature extractor
To ensure a robust feature representation for regions of varying scales, we construct a multi-level image feature pyramid Lin et al. (2017) by selecting four layers from the CLIP vision encoder and fusing them with five lightweight scale shuffle modules Zhang et al. (2023d). These layers are located at the second-to-last, fifth-to-last, eighth-to-last, and eleventh-to-last positions, respectively. Additionally, we incorporate feature coordinates Liu et al. (2018a); Wang et al. (2020) for each level to address the problem of translation invariance in CNNs. This helps make the model sensitive to absolute position information, such as the description “girl on left” in Figure 5. Finally, we use RoIAlign to extract region-level features with an output size of 14$\times$14 He et al. (2017), which ensures that sufficient detailed information is preserved. Moreover, all four level features are involved in the RoIAlign operation and fused into a single embedding as the representation of the region of interest (RoI) Liu et al. (2018b).
A more detailed framework of our GPT4RoI
1. As shown in Fig 4, We preprocess the input text by adding prefixes to retain both image information and pure text references for each region.
2. Next, we tokenize and embed the text. The image feature and region features will replace the placeholders <image> and <region{i}> respectively.
3. The resulting interleaved sequence of region & image features and language embeddings is then fed into a large language model (LLM) for further processing.
A.2 Training Details
Dialogue model The dialogue model in the demo is trained on 8 GPUs, each with 80G of memory. During the first training stage, a learning rate of 2e-5 is used with a cosine learning schedule. The batch size is 16 for 2 epochs, with a warm-up iteration set to 3000 and a warm-up ratio of 0.003. The weight decay for all modules was set to 0. During the second training stage, the learning rate is reduced to 2e-5 and the model is trained for 1 epoch. To enable end-to-end fine-tuning of the model, which includes a 7B Vicuna, Fully Sharded Data Parallel (FSDP) is enabled in PyTorch to save memory.
Downstream tasks We finetune on three datasets with different learning schedules and task prompts (as shown in Table 8). For the region caption task on Visual Genome Krishna et al. (2017), we perform fine-tuning for 4 epochs with a learning rate of 2e-5. As for Visual-7W Zhu et al. (2016), we observe that it requires a smaller learning rate of 1e-6 to stabilize the training, which is also trained in 2 epochs. On the Visual Commonsense Reasoning Zellers et al. (2019a), we fine-tune the model for 1 epoch using a learning rate of 2e-5.
Instruction of three downstream tasks. The instructions for three downstream tasks are provided in Table 8.
@@ A.2 Training Details
Region Reasoning Task on Visual-7W
### Question: <region1>,<region2>,<region3>,<region4> refers to specific areas within the photo along with their respective identifiers. I need you to answer the question. Questions are multiple-choice; you only need to pick the correct answer from the given options (A), (B), (C), or (D).
Which is the black machine under the sign?
### Answer: (A)
@@
Table 8: Task prompt of three downstream tasks.
Instruction of Single-Region Caption The instructions for single-region caption are provided in Table 9. We randomly select one as the question in training.
Instruction of Multi-Region Caption The instructions for multi-region caption are provided in Table 10. We randomly select one as the question in training.
GPT4RoI after stage1 training As shown in Figure 5, GPT4RoI is capable of identifying the category,simple attribute and position of the region.
A.3 VCR
Introduction to the VCR Dataset
@@ A.3 VCR
Preprocess of VCR
To construct a sequence of questions, we convert the explanation to a follow-up question and format them into a two-round conversation. Table 11 shows an example of the follow-up question that asks for the reasoning behind the answer.
The VCR dataset is valued for its diverse question-answer pairs that require referencing from prior question-answers to perform reasoning. Therefore, it’s crucial to assign a reference to each region in the dataset. We accomplish this by starting each conversation with a reference to all regions, e.g., There are <region1>, <region2>… in the image. This approach explicitly references every region, avoiding confusion in future analyses. Additionally, we substitute the corresponding <region{i}> in the answer with category_name at region{i} to ensure a plain text output sequence.
A.4 Other Region Reference Format
The key distinction lies in whether to incorporate the detection function into the LLM. For the method that uses textual coordinates or extra position tokens as region reference, they have to solve the following challenge:
Aligning a large number of position tokens with their corresponding positions in the image by training on a large set of datasets. But this is actually a simple rule that can be naturally implemented with the operation in detection architectures.
Modeling geometric properties can be challenging. For example, if the ground truth box is $<x_{1}=0,y_{1}=0,x_{2}=5,y_{2}=5>$, a predicted box of $<x_{1}=1,y_{1}=1,x_{2}=4,y_{2}=4>$ would be considered a better result than $<x_{1}=1,y_{1}=1,x_{2}=8,y_{2}=8>$. because it has a higher overlap with the ground truth. However, incorporating this geometric property into the next token prediction task using cross-entropy loss can be challenging. On the other hand, utilizing traditional loss functions such as L1 or IoU loss can naturally handle this geometric constraint.
Dense to Sparse  Ren et al. (2015); Zhang et al. (2023d) is a crucial design for detection performance, but embedding such an idea into the sequential form of LLM is challenging. We provide two pieces of evidence to support our argument
1. The performance of pix2seq Chen et al. (2021, 2022), which utilizes object365 Shao et al. (2019) pretrain, falls significantly behind the corresponding specialist  Zhang et al. (2022); Li et al. (2023a).
2. Even with scaled-up data and parameters, GPT4V still faces challenges in object counting Yang et al. (2023b). However, this is a trivial task for detection methods.
Another approach is to use an external detector to find the potential region of interest, whereas LLM only focuses on analyzing the corresponding region of interest. This is the motivation of GPT4RoI. It requires much less data and allows for quick adaptation to specific domain problems with the corresponding detector. However, the drawback is that the framework may appear less elegant and it assumes input contains all regions of interest that need to be analyzed.
Both approaches have their advantages and disadvantages, and academic research in both directions is thriving (including concurrent works or follow-ups on GPT4RoI). For the first approach, relevant references include Zhao et al. (2023); Chen et al. (2023b), while for the second approach, there are Pi et al. (2024); Chen et al. (2023a) besides GPT4RoI. Additionally, there has been research that explores a fusion of the two approaches, as shown in references You et al. (2023); Rasheed et al. (2023); Zhang et al. (2023b).
A.5 Failure Case Analysis
Due to limited data and instructions, GPT4RoI may fail in several landmark scenarios. We have conducted a thorough analysis and look forward to improving these limitations in future versions.
Figure 6: GPT4RoI on instruction obfuscation.
Instruction obfuscation As shown in Figure 6.(a), our multiple-region reasoning capability mainly relies on VCR, where we often use sentences that declare <region1>, <region2>, etc. at the beginning of the question. However, when users adopt the less common sentence structure to refer to regions, it can often be confused with region captions that have the highest proportion in the dataset. As shown in Figure 6.(b), because our data and instructions are mainly generated by rules, our training data does not include content with the "respectively" instruction in multi-region scenarios. This can be resolved by adding specific instructions. In future versions, we aim to develop more diverse instructions, while ensuring data balance.
Figure 7: GPT4RoI on Misidentification of fine-grained information.
Misidentification of fine-grained information within in region Although GPT4RoI has improved the fine-grained perception ability of images compared to image-level vision language models, the limited amount of region-level data results in insufficient fine-grained alignment within regions. For example, in Figure 7.(a), the model incorrectly identifies the color of the helmet, and in Figure 7.(b), it misidentifies the object in the girl’s hand. Both cases generate the corresponding answers based on the most prominent feature within the region. Using semi-supervised methods to create more region-level data may address this issue.
1. Can you provide me with a detailed description of the region in the picture marked by <region1>?
2. I’m curious about the region represented by <region1> in the picture. Could you describe it in detail?
3. What can you tell me about the region indicated by <region1> in the image?
4. I’d like to know more about the area in the photo labeled <region1>. Can you give me a detailed description?
5. Could you describe the region shown as <region1> in the picture in great detail?
6. What details can you give me about the region outlined by <region1> in the photo?
7. Please provide me with a comprehensive description of the region marked with <region1> in the image.
8. Can you give me a detailed account of the region labeled as <region1> in the picture?
9. I’m interested in learning more about the region represented by <region1> in the photo. Can you describe it in detail?
10. What is the region outlined by <region1> in the picture like? Could you give me a detailed description, please?
11. Can you provide me with a detailed description of the region in the picture marked by <region1>, please?
12. I’m curious about the region represented by <region1> in the picture. Could you describe it in detail, please?
13. What can you tell me about the region indicated by <region1> in the image, exactly?
14. I’d like to know more about the area in the photo labeled <region1>, please. Can you give me a detailed description?
15. Could you describe the region shown as <region1> in the picture in great detail, please?
16. What details can you give me about the region outlined by <region1> in the photo, please?
17. Please provide me with a comprehensive description of the region marked with <region1> in the image, please.
18. Can you give me a detailed account of the region labeled as <region1> in the picture, please?
19. I’m interested in learning more about the region represented by <region1> in the photo. Can you describe it in detail, please?
20. What is the region outlined by <region1> in the picture like, please? Could you give me a detailed description?
Table 9: A list of instructions for single-region caption.
1. Could you please give me a detailed description of these areas [<region1>, <region2>, …]?
2. Can you provide a thorough description of the regions [<region1>, <region2>, …] in this image?
3. Please describe in detail the contents of the boxed areas [<region1>, <region2>, …].
4. Could you give a comprehensive explanation of what can be found within [<region1>, <region2>, …] in the picture?
5. Could you give me an elaborate explanation of the [<region1>, <region2>, …] regions in this picture?
6. Can you provide a comprehensive description of the areas identified by [<region1>, <region2>, …] in this photo?
7. Help me understand the specific locations labeled [<region1>, <region2>, …] in this picture in detail, please.
8. What is the detailed information about the areas marked by [<region1>, <region2>, …] in this image?
9. Could you provide me with a detailed analysis of the regions designated [<region1>, <region2>, …] in this photo?
10. What are the specific features of the areas marked [<region1>, <region2>, …] in this picture that you can describe in detail?
11. Could you elaborate on the regions identified by [<region1>, <region2>, …] in this image?
12. What can you tell me about the areas labeled [<region1>, <region2>, …] in this picture?
13. Can you provide a thorough analysis of the specific locations designated [<region1>, <region2>, …] in this photo?
14. I am interested in learning more about the regions marked [<region1>, <region2>, …] in this image. Can you provide me with more information?
15. Could you please provide a detailed description of the areas identified by [<region1>, <region2>, …] in this photo?
16. What is the significance of the regions labeled [<region1>, <region2>, …] in this picture?
17. I would like to know more about the specific locations designated [<region1>, <region2>, …] in this image. Can you provide me with more information?
18. Can you provide a detailed breakdown of the regions marked [<region1>, <region2>, …] in this photo?
19. What specific features can you tell me about the areas identified by [<region1>, <region2>, …] in this picture?
20. Could you please provide a comprehensive explanation of the locations labeled [<region1>, <region2>, …] in this image?
Table 10: A list of instructions for multiple-region caption.
1. Why? 2. What’s the rationale for your decision 3. What led you to that conclusion? 4. What’s the reasoning behind your opinion? 5. Can you explain the basis for your thinking? 6. What factors influenced your perspective? 7. How did you arrive at that perspective? 8. What evidence supports your viewpoint? 9. What’s the logic behind your argument? 10. Can you provide some context for your opinion? 11. What’s the basis for your assertion? 12. What experiences have shaped your perspective? 13. What assumptions underlie your reasoning? 14. What’s the foundation of your assertion? 15. What’s the source of your reasoning? 16. What’s the motivation behind your decision? 17. What’s the impetus for your belief? 18. What’s the driving force behind your conclusion? 19. What’s your reasoning? 20. What makes you say that? 21. What’s the story behind that? 22. What’s your thought process? 23. What’s the deal with that? 24. What’s the logic behind it? 25. What’s the real deal here? 26. What’s the reason behind it? 27. What’s the rationale for your opinion? 28. What’s the background to that? 29. What’s the evidence that supports your view? 30. What’s the explanation for that?
Table 11: A list of instructions for the second round chat in VCR.

### v5

@@ 5.2 Region Recognization
We also conduct evaluations on several specific region-level understanding tasks to demonstrate the superior performance of GPT4RoI. The instruction template for these tasks is provided in Table 5.
Table 5: Task prompt of three downstream tasks.
@@
Open-Vocabulary Recognition The results are obtained by calculating the semantic similarity between the generated region caption of GPT4RoI and the vocabulary lists of each dataset and then selecting the category with the highest similarity as the final result (as the study [86]). The input to the CLIP-Surgery-ViT-L is the cropped region with a size of 512$\times$512. As shown in Table 6, constrained by the resolution of 224$\times$224, GPT4RoI exhibits slightly lower performance than CLIP-Surgery-ViT-L only at the instance level. Considering all metrics score on two datasets, GPT4RoI demonstrates robust and comprehensive region recognition capabilities compared to our concurrent works that employ alternative reference formats.
Table 6: Recognition performance on panoptic segmentation (PQ), instance segmentation (AP) and semantic segmentation (mIoU) upon the validation sets of Cityscapes [14] and ADE20K [98]. means concurrent work.
Method Cityscapes ADE20K-150
PQ AP mIoU PQ AP mIoU
CLIP-Surgery-ViT-L 27.24 28.35 21.92 26.55 29.70 21.42
Kosmos-2 12.09 9.81 13.71 6.53 4.33 5.40
Shikra-7B 17.80 11.53 17.77 27.52 20.35 18.24
GPT4RoI-7B 34.70 21.93 36.73 36.32 26.08 25.82
Model BLEU@4 METEOR ROUGE CIDEr
GRiT - 17.1 - 142.0
Shikra-7B 8.9 15.2 30.3 115.8
GPT4RoI-7B 10.5 16.5 33.4 134.5
GPT4RoI-7B 11.5 17.4 35.0 145.2
GPT4RoI-13B 11.7 17.6 35.2 146.8
Model Code #Params Val Acc.(%) Test Acc.(%)
ViLBERT [43] Y 221M 72.4 74.5 54.0 73.3 74.6 54.8
Unicoder-VL [30] Y - 72.6 74.5 54.5 73.4 74.4 54.9
VLBERT-L [63] Y 383M 75.5 77.9 58.9 75.8 78.4 59.7
UNITER-L [10] Y 303M - - - 77.3 80.8 62.8
ERNIE-ViL-L [84] Y - 78.52 83.37 65.81 79.2 83.5 66.3
MERLOT [91] Y 223M - - - 80.6 80.4 65.1
VILLA-L [20] Y - 78.45 82.57 65.18 78.9 82.8 65.7
RESERVE-L [90] Y 644M - - - 84.0 84.9 72.0
VQA-GNN-L [74] Y 1B+ - - - 85.2 86.6 74.0
GPT4RoI-7B (ours) Y 7B+ 87.4 89.6 78.6 - - -
VLUA+@Kuaishou N - - - - 84.8 87.0 74.0
KS-MGSR@KDDI Research and SNAP N - - - - 85.3 86.9 74.3
SP-VCR@Shopee N - - - - 83.6 88.6 74.4
HunYuan-VCR@Tencent N - - - - 85.8 88.0 75.6
Human Performance [88] - - - - - 91.0 93.0 85.0
GPT4RoI-13B (ours) Y 13B+ - - - 89.4 91.0 81.6
Region Caption We report the scores of BLEU, METEOR, ROUGE, and CIDEr for different methods on the validation set of Visual Genome [29]. As shown in Table 7, compared to the concurrent work Shikra [8] with textual coordinates as region reference, GPT4RoI shows clearly better region caption ability. After fine-tuning, GPT4RoI outperforms the previous state-of-the-art specialist model GRiT [76] by a significant margin, without any additional techniques or tricks. Additionally, we observe that the performance of GPT4RoI-7B and GPT4RoI-13B is comparable, suggesting that the bottleneck in performance lies in the design of the visual module and the availability of region-text pair data. These areas can be explored further in future work.
5.3 Region Reasoning
Visual Commonsense Reasoning [88] employs multiple-choice settings that require both recognition ability and commonsense reasoning to select the correct choice. Calculating the similarity of LLM’s output with the correct choice is quite a challenge. This is because it is hard to find a model that can capture the reasoning similarity. Therefore, we finetune GPT4RoI to align with the answer format, following conventional methods.
@@ 5.3 Region Reasoning
6 Conclusions
7 Acknowledgement
This project is supported by the National Key R&D Program of China
No.2022ZD0161000.